FEMTO-FANTASIO: A VERSATILE EXPERIMENTAL SET-UP TO INVESTIGATE MOLECULAR COM-PLEXES

\underline{K. DIDRICHE}, C. LAUZIN, X. DE GHELLINCK, T. FOLDES and M. HERMAN, Service de Chimie quantique et Photophysique CP160/09, Faculté des Sciences, Université Libre de Bruxelles (U.L.B.), Av. Roosevelt, 50, B-1050, Bruxelles, Belgium.

Several improvements have been made on the apparatus developed in Brussels for the study of jet-cooled molecules, named FANTASIO ^{a}^{b}^{c}, for ”Fourier trANsform, Tunable, diode and quadrupole mAss spectrometers interfaced to a Supersonic expansIOn”. The upgraded setup, called Femto-FANTASIO, is presented. These improvements include: a doubling of the pumping efficiency, by adding another turbomolecular pump Leybold MAG-W3200 CT; an increase in sensitivity in the 1.5 μm range thanks to a new CRDS system, consisting in new cavity mirrors, leading to a ring down time of 125 μs and a new detector, decreasing the noise level; a new injection system to probe samples liquid at STP conditions; a temperature controlled nozzle/slit; and as an alternative probe technique, an absorption source, tunable from 3000 to 9000 cm^{−1}, coupling an optical parametric oscillator (OPO), pumped by a Ti:Sa femtosecond laser, to a high resolution continuous scan Fourier transform interferometer. Femto FANTASIO will be used to investigate molecular complexes. First results and achievements of this new setup are presented.

aM. Herman, K. Didriche, D. Hurtmans, B. Kizil, P. Macko, A. Rizopoulos and P. Van Poucke Mol. Phys. \underline{105}(5-7), 815 (2007).

bK. Didriche, C. Lauzin, P. Macko, W. J. Lafferty, R. J. Saykally and M. Herman Chem. Phys. Lett. \underline{463}, 345 (2008).

cC. Lauzin, K. Didriche, P. Macko, J. Demaison, J. Liévin and M. Herman J. Phys. Chem. A \underline{113}, 2359 (2009).